The traffic light turns green and the driver of a high performance car slams the accelerator to the floor. The accelerator registers 22.0m/s2.Convert this reading to km/min2,
1
Expert's answer
2018-11-26T16:07:09-0500

22.0 m/s^2 =22.0 ( 1/1000 km)/((1/60)^2 min^2 )=79.2 km/(min^2 )